• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armando Acha

    Log entry time 00:05:41 on May30,2007

    Entry number 203545

    keyword=Swing Shift Summary

    Swing Shift 5/29/2007

    SL: Armando Acha

    Shift Workers:

    Mina Kat
    Javier Gomez
    Petratos Katramatou

    No beam until 7:30 pm
    Target warm up at the beggining of the shift to go to He4. This caused some shifting in the target apparatus position. This required controlled access which started in previous shift. It was finished at 5:00pm approx.

    Starting at 7:30pm we dedicated to locate the new coordinates of beam positions using BeO target to see where the beam is with the camera and Carbon hole target to figure out the new coordinates now to be:

    y=2.0 for both A and B

    process completed at 21:09 see log entry 203537.

    Runs 3829 - 3842 are test runs on BeO and C hole.

    Back to BeO target:

    After that we requested harp scans to make sure the size of the beam spot is correct. X sigma was too small 0.05 we requested it to be increased to 0.1 at least. See Elog entry #1378726.

    22:10 new harp scan X sigma is good but now Y sigma needs to be fixed. We expect new change and harp scan to get the right beam spot size.

    22:15 Configuration change in the spectrometers to run at Q2=8 fm-2 on He4. see halog entry 203541.

    22:35 harp scan now looks fine see Elog entry 1378731. We will verify everything is ok , beam, raster , spot size before moving to He4.

    23:47 MCC called to advice us of doing a new harp scan since the one they performed was not that correct one. New result .96 and .94 for the X and Y sigmas. We will run with these ones. Elog entry 1 378744.

    23:56 Requested raster to be turn on.

    0:00 Finally everything ok to start taking data production in He 4.